Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find

Upper P 80

0.30.3

0.40.4

0.40.4

0.40.4

0.40.4

0.40.4

0.50.5

0.60.6

0.60.6

0.60.6

0.70.7

0.80.8

0.80.8

0.80.8

0.90.9

0.90.9

1.21.2

1.51.5

1.81.8

1.81.8

1.91.9

2.22.2

2.22.2

2.32.3

2.42.4

2.52.5

2.62.6

2.92.9

3.23.2

3.93.9

4.34.3

4.44.4

5.55.5

5.65.6

6.96.9

7.17.1

8.68.6

8.98.9

10.410.4

10.510.5

10.510.5

11.511.5

11.611.6

12.112.1

12.212.2

13.213.2

13.713.7

14.714.7

15.315.3

25.525.5

Upper P 80= ?Mbps

​(Type an integer or a decimal. Do not​ round.)
Answer #1

Answer

P80 means 80th percentile

percentile formula is

percentile = n*(percentile/100)

we have n = 50 (total data values) and percentile = 80

so, we get

P80 = 50*(80/100)

= 50*0.8

= 40th score

using the data, we get 40th score = 10.5

so, P80 = 10.5